# A Dynamic Recurrent Adjacency Memory Network for Mixed-Generation Power System Stability Forecasting

### v1

@@ I Introduction
The assessment and forecasting of stability in modern power systems, characterized by high penetration levels of inverter-based resources (IBRs), presents growing challenges [1, 2, 3]. Traditional stability assessment methods, such as eigenvalue analysis for small-signal conditions and time-domain simulations for transient behavior, are based on offline studies of fixed system models. These methods, though reliable under conventional setups, encounter limitations in scalability and accuracy as the system operating point evolves more rapidly due to the variability introduced by IBRs [4]. As the share of IBRs increases, the dynamic interactions between synchronous generators (SGs), grid-forming (GFM), and grid-following (GFL) converters become more complex, causing significant variation in stability margins and transient response characteristics. These evolving dynamics challenge the assumptions underlying classical techniques, motivating the exploration of data-driven methods as a promising alternative. Such approaches aim to deliver fast and generalizable stability forecasts using real-time measurements, thereby reducing reliance on computationally intensive simulations and fixed system models [5, 2].
@@ I-A Related Work
[6, 7] Modal / time-domain analysis Foundational baselines
[8] GNNs Stability risk assessment
[9] GNNs Large-scale analysis
[10] GraphSAGE-A Multi-task transient stability assessment
[11] Swin Transformer Transient stability assessment
@@ I-A1 Classical Methods for Stability Assessment
Small-signal and transient stability in power systems are traditionally assessed using eigenvalue-based modal analysis and time-domain simulations. Modal analysis enables the study of system dynamics by linearizing the system around a steady-state operating point [6]. It is effective for tuning power system stabilizers and evaluating oscillatory behavior under small disturbances. However, its reliance on accurate linear models and fixed operating points limits its adaptability to fast-changing system conditions. Time-domain simulations numerically solve differential algebraic equations to capture the full nonlinear response under large disturbances. While highly accurate, especially for transient stability analysis, these approaches are computationally intensive and unsuitable for real-time applications [7]. Both fundamental techniques struggle to scale under high variability and limited observability conditions induced by widespread IBR deployment.
@@ I-A2 Data-Driven Machine Learning Models
To overcome limitations in classical stability assessment methods, data-driven techniques have been increasingly adopted for stability classification and prediction. Classical ML models such as support vector machines, decision trees, random forests, and XGBoost have been applied using static features like pre-fault voltages, rotor angles, and bus frequencies [12]. While these methods are lightweight and easy to interpret, they often rely on manual feature engineering and fail to capture temporal dependencies inherent in system dynamics.
Recurrent models such as long short-term memory recurrent neural networks (LSTM-RNNs) and gated recurrent units have shown promise for capturing sequential patterns in phasor measurement unit (PMU) data [12, 5]. These approaches improve temporal awareness but typically treat each system node independently, ignoring the spatial and structural dependencies of the grid. As a result, they may overfit to specific scenarios and exhibit poor generalization under unseen operating conditions.
@@ I-A3 Graph Neural Networks in Power Systems
Graph neural networks (GNNs) provide a natural framework for modeling power systems, where buses and lines form an inherent graph structure. Early GNN applications in this domain use graph convolutional networks (GCNs) or graph attention networks (GATs) to incorporate topological information into load forecasting, fault localization, or stability classification tasks [8, 9]. However, most of these models use static adjacency matrices derived from physical bus connectivity or line admittance, which do not capture dynamic electrical interactions or modal coupling effects that vary with generation mix, load level, or fault type.
@@ I-A4 Physics-Informed and Hybrid Approaches
To address the gap between physical fidelity and data-driven generalization, hybrid models have been developed to incorporate physics-based features or constraints into learning frameworks. Some studies include inertia constants, damping coefficients, or participation factors as input features to neural networks [13, 14], while others enforce physical laws via regularization terms in the loss function [15]. These models improve interpretability and generalization but often require extensive modeling effort or access to detailed system parameters.
A promising direction involves the use of dynamic mode decomposition (DMD) to extract dominant spatiotemporal modes from measurement data. DMD has been applied for post-event oscillation analysis, inertia estimation, and Koopman operator approximation in power systems [16]. Variants such as time-delayed DMD (TD-DMD) [17], kernel extended DMD (KE-DMD) [18], and Hankel DMD [19] improve robustness under noise and nonlinearity. However, most DMD applications remain limited to offline analysis and are not integrated with graph-based learning or used as inputs to real-time forecasting architectures. The framework proposed in this study overcomes these limitations by integrating DMD-derived spectral interactions directly within recurrent graph convolutions, enabling real-time learning of evolving modal relationships. In contrast to previous DMD or GNN methods, it jointly models temporal and spatial dependencies through dynamic adjacency updates, providing unified and interpretable stability forecasting across varying operating conditions.
@@ II-A Sliding-Window DMD for Adjacency Matrix Extraction
This section outlines the construction of multi-layer adjacency matrices from raw time-series data using a sliding-window DMD framework. DMD is a data-driven technique that approximates system dynamics by decomposing time-series snapshots into spatial modes associated with growth or decay rates and oscillatory frequencies [16]. In this work, DMD is applied over sliding windows to capture the evolving spectral content of the system, thereby reflecting transient variations in modal interactions. For each window, the decomposition is truncated to rank $r=5$, which retains the dominant electromechanical and converter-driven modes typically governing stability, while filtering out noise and reducing computational complexity. The resulting modes form a five-layer adjacency tensor, with each layer encoding a distinct dynamic property of the grid.
A linear operator $A$ is estimated via DMD to characterize the temporal evolution of discrete-time system states $\mathbf{x}_{k+1}\approx A\,\mathbf{x}_{k}$, where $A\in\mathbb{R}^{n\times n}$ and $\mathbf{x}_{k}\in\mathbb{R}^{n}$ is the state at sample $k$. For a window of $T$ samples with sampling interval $\Delta t$, the snapshot matrices $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$ are formed in (1)-(2):
@@ III-A A Ternary Parameterization System
The penetration levels of SG, GFM, and GFL components within an electrical grid can be effectively illustrated using barycentric coordinates on a ternary grid [5, 20]. Each grid point in this system corresponds to a distinct convex combination of penetration levels which sums exactly to 100 (%), representing the total dispatch in the system.
@@ III-B1 A Modified 9-Bus Model
A modified version of the IEEE standard 9-bus system [5] is used as a representative benchmark for assessing the proposed stability prediction framework. In this model, two of the synchronous generator units are each replaced with a GFM and a GFL inverter. In this model, the SG is equipped with a governor, excitation control, and a power system stabilizer. The GFM inverter is droop-control-based, and the GFL includes DC-side dynamics.
@@ III-B2 HVDC Model
The U-shaped multi-terminal HVDC system represents a more complex mixed-generation grid comprising GFL renewable energy resources, BESS, SG-based hydro storage, and a strong grid interconnection distributed across four different AC areas interconnected through a four-terminal HVDC network, as shown in Fig. 2 [5]. The system comprises two SGs equipped with the same controllers as in the modified 9-bus model, representing pumped hydro (PH) units. The GFM battery storage units are placed in the same areas as the GFL PV generation, which are modeled as in the modified 9-bus model.
@@ III-B3 A Modified 39-Bus System
A modified IEEE 39-bus system [21] is also included to assess scalability. Generators 3, 8, and 9 are replaced by GFM units, and generators 4, 7, and 10 are replaced by GFL units. The dataset generation follows the same procedure described above, with each operating point simulated for 60 s. Unlike the 9-bus and HVDC systems, only unperturbed operating scenarios are simulated for the 39-bus system. This design choice is intended to demonstrate the scalability of the proposed framework under larger network sizes, while limiting the dataset complexity to steady-state variations in the generation mix.
@@ III-C Data Labeling and Preparation
Since the objective of DRAMN is to predict system stability based on current time-series data, all sliding windows extracted from measurements corresponding to a given generation mix scenario are assigned a common stability label. This label is assigned by the modal analysis function of PowerFactory [22], which computes eigenvalues at the end of each simulation. A case is classified as “1” (unstable) if it meets any of the following criteria:
@@
    Exhibits damping ratios $\zeta<3\%$ [25]
@@ IV-B Model Selection and Benchmarking
ST-AGNet [26] uses spatial-temporal attention and adaptive graph structure learning to capture complex dependencies in power networks. For this study, predefined electrical connections are used to build graph topology, while temporal features are extracted via convolutional layers. The attention mechanism operates across both spatial and temporal dimensions to model dynamic system relationships. The CNN+GAT [10] model integrates a CNN-based encoder with graph attention networks to classify stability and identify critical generators in a multi-task setup. Buses are mapped to nodes and branches to edges, with CNN handling temporal variation and GAT extracting spatial features. The input tensor is reshaped for temporal coherence, and the task is modified for binary classification. The transformer-based TSA model [11] employs an encoder-decoder architecture with attention to assess transient stability. Transformer blocks include feed-forward networks and normalization layers. Positional encoding captures sequence structure, while attention mechanisms model temporal dependencies across multiple scales. The PowerGNN model [27] constructs graph representations via DMD and applies a graph over-sampling method to address power-law node distributions. Multiple adjacency layers—based on participation factors, mode characteristics, and eigenvalues—are normalized and used for graph convolution, capturing both topological and dynamic system behavior.
The DMD algorithm is also replaced with three advanced variants to evaluate the optimal method for constructing adjacency matrices, as shown in Table IV. Neural DMD (N-DMD) [28] integrates neural network-based autoencoders with spectral decomposition, where observations are encoded into a lifted space and forecasts are made through DMD analysis of the encoded data. Kernel extended DMD (KE-DMD) [29] leverages radial basis function kernels to capture nonlinear dynamics without explicit feature mappings. Time-Delayed DMD (TD-DMD) [30] embeds temporal delays into the state representation using parameters selected via mutual information analysis, enriching the spectral structure for stability assessment.
@@ 19 Nodes
Given the five-layer structure of each adjacency matrix, five corresponding graph networks are instantiated per time step to capture multi-dimensional connectivity. The cumulative node strength $W_{n}$ for the $n^{th}$ node of each graph in each of the five layers is defined as $W_{n}=\sum_{t=19000}^{30000}\sum_{j=1}^{N}A^{(l)}_{nj}(t)$, where $A_{nj}^{(l)}(t)$ represents the edge weight between nodes $n$ and $j$ in layer $l$ at time $t$, and $N$ denotes the total number of nodes in the network. To ensure scale-invariant node selection, the raw cumulative node strengths in each layer are normalized using min-max normalization, such that $\tilde{W}_{n}=\frac{W_{n}-\min(W)}{\max(W)-\min(W)}$, where $\tilde{W}_{n}$ represents the normalized strength of node $n$. These node strengths are then summed to produce an aggregated connectivity graph, defined as $\tilde{W}n^{(composite)}=\sum_{l=1}^{5}\tilde{W}_{n}^{(l)}$.

### v2

@@ I Introduction
The assessment and forecasting of stability in modern power systems, characterized by high penetration levels of inverter-based resources (IBRs), presents growing challenges [Hatziargyriou2021, alimi2020review, Milano2018]. Traditional stability assessment methods, such as eigenvalue analysis for small-signal conditions and time-domain simulations for transient behavior, are based on offline studies of fixed system models. These methods, though reliable under conventional setups, encounter limitations in scalability and accuracy as the system operating point evolves more rapidly due to the variability introduced by IBRs [zhou2024efficient]. As the share of IBRs increases, the dynamic interactions between synchronous generators (SGs), grid-forming (GFM), and grid-following (GFL) converters become more complex, causing significant variation in stability margins and transient response characteristics. These evolving dynamics challenge the assumptions underlying classical techniques, motivating the exploration of data-driven methods as a promising alternative. Such approaches aim to deliver fast and generalizable stability forecasts using real-time measurements, thereby reducing reliance on computationally intensive simulations and fixed system models [bertozzi2024data, alimi2020review].
@@ I-A Related Work
[kelada2025revisiting, lara2023revisiting] Modal / time-domain analysis Foundational baselines
[chen2024real] GNNs Stability risk assessment
[zhu2024scaling] GNNs Large-scale analysis
[huang2024multi] GraphSAGE-A Multi-task transient stability assessment
[zhao2024swin] Swin Transformer Transient stability assessment
@@ I-A1 Classical Methods for Stability Assessment
Small-signal and transient stability in power systems are traditionally assessed using eigenvalue-based modal analysis and time-domain simulations. Modal analysis enables the study of system dynamics by linearizing the system around a steady-state operating point [kelada2025revisiting]. It is effective for tuning power system stabilizers and evaluating oscillatory behavior under small disturbances. However, its reliance on accurate linear models and fixed operating points limits its adaptability to fast-changing system conditions. Time-domain simulations numerically solve differential algebraic equations to capture the full nonlinear response under large disturbances. While highly accurate, especially for transient stability analysis, these approaches are computationally intensive and unsuitable for real-time applications [lara2023revisiting]. Both fundamental techniques struggle to scale under high variability and limited observability conditions induced by widespread IBR deployment.
@@ I-A2 Data-Driven Machine Learning Models
To overcome limitations in classical stability assessment methods, data-driven techniques have been increasingly adopted for stability classification and prediction. Classical ML models such as support vector machines, decision trees, random forests, and XGBoost have been applied using static features like pre-fault voltages, rotor angles, and bus frequencies [alimi2021review]. While these methods are lightweight and easy to interpret, they often rely on manual feature engineering and fail to capture temporal dependencies inherent in system dynamics.
Recurrent models such as long short-term memory recurrent neural networks (LSTM-RNNs) and gated recurrent units have shown promise for capturing sequential patterns in phasor measurement unit (PMU) data [alimi2021review, bertozzi2024data]. These approaches improve temporal awareness but typically treat each system node independently, ignoring the spatial and structural dependencies of the grid. As a result, they may overfit to specific scenarios and exhibit poor generalization under unseen operating conditions.
@@ I-A3 Graph Neural Networks in Power Systems
Graph neural networks (GNNs) provide a natural framework for modeling power systems, where buses and lines form an inherent graph structure. Early GNN applications in this domain use graph convolutional networks (GCNs) or graph attention networks (GATs) to incorporate topological information into load forecasting, fault localization, or stability classification tasks [chen2024real, zhu2024scaling]. However, most of these models use static adjacency matrices derived from physical bus connectivity or line admittance, which do not capture dynamic electrical interactions or modal coupling effects that vary with generation mix, load level, or fault type.
@@ I-A4 Physics-Informed and Hybrid Approaches
To address the gap between physical fidelity and data-driven generalization, hybrid models have been developed to incorporate physics-based features or constraints into learning frameworks. Some studies include inertia constants, damping coefficients, or participation factors as input features to neural networks [linaro2023continuous, cao2023physics], while others enforce physical laws via regularization terms in the loss function [falas2025robust]. These models improve interpretability and generalization but often require extensive modeling effort or access to detailed system parameters.
A promising direction involves the use of dynamic mode decomposition (DMD) to extract dominant spatiotemporal modes from measurement data. DMD has been applied for post-event oscillation analysis, inertia estimation, and Koopman operator approximation in power systems [korda2018convergence]. Variants such as time-delayed DMD (TD-DMD) [nedzhibov2025blind], kernel extended DMD (KE-DMD) [philipp2025error], and Hankel DMD [sakib2025learning] improve robustness under noise and nonlinearity. However, most DMD applications remain limited to offline analysis and are not integrated with graph-based learning or used as inputs to real-time forecasting architectures. The framework proposed in this study overcomes these limitations by integrating DMD-derived spectral interactions directly within recurrent graph convolutions, enabling real-time learning of evolving modal relationships. In contrast to previous DMD or GNN methods, it jointly models temporal and spatial dependencies through dynamic adjacency updates, providing unified and interpretable stability forecasting across varying operating conditions.
@@ II-A Sliding-Window DMD for Adjacency Matrix Extraction
This section outlines the construction of multi-layer adjacency matrices from raw time-series data using a sliding-window DMD framework. DMD is a data-driven technique that approximates system dynamics by decomposing time-series snapshots into spatial modes associated with growth or decay rates and oscillatory frequencies [korda2018convergence]. In this work, DMD is applied over sliding windows to capture the evolving spectral content of the system, thereby reflecting transient variations in modal interactions. For each window, the decomposition is truncated to rank $r=5$, which retains the dominant electromechanical and converter-driven modes typically governing stability, while filtering out noise and reducing computational complexity. The resulting modes form a five-layer adjacency tensor, with each layer encoding a distinct dynamic property of the grid.
A linear operator AA is estimated via DMD to characterize the temporal evolution of discrete-time system states $\mathbf{x}_{k+1}\approx A\,\mathbf{x}_{k}$, where $A\in\mathbb{R}^{n\times n}$ and $\mathbf{x}_{k}\in\mathbb{R}^{n}$ is the state at sample $k$. For a window of $T$ samples with sampling interval $\Delta t$, the snapshot matrices $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$ are formed in (1)-(2):
@@ III-A A Ternary Parameterization System
The penetration levels of SG, GFM, and GFL components within an electrical grid can be effectively illustrated using barycentric coordinates on a ternary grid [bertozzi2024data, BERTOZZI2025planning]. Each grid point in this system corresponds to a distinct convex combination of penetration levels which sums exactly to 100 (%), representing the total dispatch in the system.
@@ III-B1 A Modified 9-Bus Model
A modified version of the IEEE standard 9-bus system [bertozzi2024data] is used as a representative benchmark for assessing the proposed stability prediction framework. In this model, two of the synchronous generator units are each replaced with a GFM and a GFL inverter. In this model, the SG is equipped with a governor, excitation control, and a power system stabilizer. The GFM inverter is droop-control-based, and the GFL includes DC-side dynamics.
@@ III-B2 HVDC Model
The U-shaped multi-terminal HVDC system represents a more complex mixed-generation grid comprising GFL renewable energy resources, BESS, SG-based hydro storage, and a strong grid interconnection distributed across four different AC areas interconnected through a four-terminal HVDC network, as shown in Fig. 2 [bertozzi2024data]. The system comprises two SGs equipped with the same controllers as in the modified 9-bus model, representing pumped hydro (PH) units. The GFM battery storage units are placed in the same areas as the GFL PV generation, which are modeled as in the modified 9-bus model.
@@ III-B3 A Modified 39-Bus System
A modified IEEE 39-bus system [athay1979] is also included to assess scalability. Generators 3, 8, and 9 are replaced by GFM units, and generators 4, 7, and 10 are replaced by GFL units. The dataset generation follows the same procedure described above, with each operating point simulated for 60 s. Unlike the 9-bus and HVDC systems, only unperturbed operating scenarios are simulated for the 39-bus system. This design choice is intended to demonstrate the scalability of the proposed framework under larger network sizes, while limiting the dataset complexity to steady-state variations in the generation mix.
@@ III-C Data Labeling and Preparation
Since the objective of DRAMN is to predict system stability based on current time-series data, all sliding windows extracted from measurements corresponding to a given generation mix scenario are assigned a common stability label. This label is assigned by the modal analysis function of PowerFactory [powerfactory_manual], which computes eigenvalues at the end of each simulation. A case is classified as “1” (unstable) if it meets any of the following criteria:
@@
    Exhibits damping ratios $\zeta<3\%$ [IEEE1110_2019]
@@ IV-B Model Selection and Benchmarking
ST-AGNet [zhang2024stagnet] uses spatial-temporal attention and adaptive graph structure learning to capture complex dependencies in power networks. For this study, predefined electrical connections are used to build graph topology, while temporal features are extracted via convolutional layers. The attention mechanism operates across both spatial and temporal dimensions to model dynamic system relationships. The CNN+GAT [huang2024multi] model integrates a CNN-based encoder with graph attention networks to classify stability and identify critical generators in a multi-task setup. Buses are mapped to nodes and branches to edges, with CNN handling temporal variation and GAT extracting spatial features. The input tensor is reshaped for temporal coherence, and the task is modified for binary classification. The transformer-based TSA model [zhao2024swin] employs an encoder-decoder architecture with attention to assess transient stability. Transformer blocks include feed-forward networks and normalization layers. Positional encoding captures sequence structure, while attention mechanisms model temporal dependencies across multiple scales. The PowerGNN model [li2025power] constructs graph representations via DMD and applies a graph over-sampling method to address power-law node distributions. Multiple adjacency layers—based on participation factors, mode characteristics, and eigenvalues—are normalized and used for graph convolution, capturing both topological and dynamic system behavior.
The DMD algorithm is also replaced with three advanced variants to evaluate the optimal method for constructing adjacency matrices, as shown in Table IV. Neural DMD (N-DMD) [iwata2023neural] integrates neural network-based autoencoders with spectral decomposition, where observations are encoded into a lifted space and forecasts are made through DMD analysis of the encoded data. Kernel extended DMD (KE-DMD) [williams2015kernel] leverages radial basis function kernels to capture nonlinear dynamics without explicit feature mappings. Time-Delayed DMD (TD-DMD) [bronstein2022spatiotemporal] embeds temporal delays into the state representation using parameters selected via mutual information analysis, enriching the spectral structure for stability assessment.
@@ 19 Nodes
Given the five-layer structure of each adjacency matrix, five corresponding graph networks are instantiated per time step to capture multi-dimensional connectivity. The cumulative node strength $W_{n}$ for the $n^{th}$ node of each graph in each of the five layers is defined as $W_{n}=\sum_{t=19000}^{30000}\sum_{j=1}^{N}A^{(l)}_{nj}(t)$, where $A_{nj}^{(l)}(t)$ represents the edge weight between nodes $n$ and $j$ in layer $l$ at time $t$, and NN denotes the total number of nodes in the network. To ensure scale-invariant node selection, the raw cumulative node strengths in each layer are normalized using min-max normalization, such that $\tilde{W}_{n}=\frac{W_{n}-\min(W)}{\max(W)-\min(W)}$, where $\tilde{W}_{n}$ represents the normalized strength of node $n$. These node strengths are then summed to produce an aggregated connectivity graph, defined as $\tilde{W}n^{(composite)}=\sum_{l=1}^{5}\tilde{W}_{n}^{(l)}$.